OpenStudy (anonymous):

What is the solution set for |2x – 3| = 17? 10 and –7 10 and –10 7 and –7 –10 and 7

OpenStudy (anonymous):

|2x-3| = 17 |2x| = 20 |x| = 10 x = 10 and -10

OpenStudy (anonymous):

wait nvm

OpenStudy (apoorvk):

if \(|f| = g\), then, \(f = \pm g\) Use this^ funda.

OpenStudy (anonymous):

If you plug in 10, the solution is 17 If you plug in -7, the solutionis 17 If you plug in -10, the solution is 23 If you plug in -4, the solution is 11

OpenStudy (anonymous):

So the solution is 10 and -7

OpenStudy (diyadiya):

2x-3 =17 2x=20 x=10 2x-3=-17 2x=-17+3 2x=-14 x=-7

OpenStudy (anonymous):

The answer should be 10 and -7, if you're talking real #'s (imaginary numbers are an entirely different topic and not what I assume you're working on). Take a look at the graph I attached as a small image file. :-) Assume you have multiple equations here, and you are finding where #1 and #2 intersect the constant value (that's the xc): \[x_c = 17\] \[x_1 = 2x-3\] \[x_2 = -2x+3\] What's cool about this method is that you can solve any in equality set, even larger polynomials or one inequality polynomial equal to another polynomial. It helps sometimes to sketch it out visually. Hope this helps!